If the first and third of three consecutive odd integers are added, the result is 69 less than five times the second integer. Find the third integer.

I dont understand how to solve this problem or even how to set problem up

Answer by mananth(16949)

let the odd integers be x,x+2,x+4
..
x+x+4 =5(x+2)-69
2x+4=5x+10-69
2x+4=5x-59
add -4 to both sides
2x+4-4=-59-4+5x
2x=-63+5x
add -5x to both sides
2x-5x= 63+5x-5x
-3x=-63
divide by -3
-3x/-3 = -63/-3
x=21
the numbers are 21, 23, 25
CHECK
21+25 = 5*(23)-69
46=115-69
46=46